OpenStudy (anonymous):

What is the product of 2.5 × 10–15 and 3.9 × 1026? 9.75 × 1041 6.4 × 101 6.4 × 1041 9.75 × 1011

OpenStudy (anonymous):

@chucky00

OpenStudy (anonymous):

yea i kinda suck at math

OpenStudy (anonymous):

haha!

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

uhhh what am I supposed to do? this is confusing.... :(

OpenStudy (anonymous):

2.5 × 10^–15 and 3.9 × 10^26?

OpenStudy (anonymous):

2.5 x 10^-15 = ? 3.9 x 10^26 = ?

OpenStudy (haleyelizabeth2017):

oh! okay one second

OpenStudy (anonymous):

lol thanks

OpenStudy (haleyelizabeth2017):

if I do that, that is none of the options....

OpenStudy (anonymous):

u have to add them i think? lol idk

OpenStudy (haleyelizabeth2017):

I'm not sure on this one....sorry!

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

\(\large (2.5\times10^{-15})(3.9\times10^{26})\) \(\large (2.5\times3.9)(10^{-15}\times10^{26})\) Can you do the rest yourself? HINT: \(\large x^{m}\times x^{n}=x^{m+n}\)

OpenStudy (anonymous):

its A!

OpenStudy (studygurl14):

Not quite

OpenStudy (studygurl14):

Remember that you are adding -15 and 26 \(\large -15+26=26-15=?\)

OpenStudy (studygurl14):

the 9.75 part is correct tho

OpenStudy (anonymous):

D! sorry i meant D

OpenStudy (studygurl14):

yea D. nice job :)